OpenStudy (anonymous):

Describe in detail how you would construct a 95% confidence interval for a set of 30 data points whose mean is 20 and population standard deviation is 3. Be sure to show that you know the formula and how to plug into it. Also, clearly state the margin of error.

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

i don't think were making one i think we are just describing how to make it since we aren't given 30 points to use

OpenStudy (anonymous):

@perl

OpenStudy (perl):

$$ \Large Confidence~ Interval = Point~ estimate \pm Margin~of~Error $$

OpenStudy (anonymous):

for margin of error i got 5.88

OpenStudy (anonymous):

wait what?

OpenStudy (perl):

how did you get 5.88 for margin of error

OpenStudy (anonymous):

critical value times standard deviation

OpenStudy (anonymous):

oh okay what is the formula

OpenStudy (perl):

one moment, checking

OpenStudy (anonymous):

ok

OpenStudy (perl):

$$ \Large{ \text{confidence interval for sample mean is: } \\ \bar x \pm z^* \frac{\sigma }{\sqrt n} \\~\\ \text{confidence interval for sample proportion is: } \\ \hat p \pm z^* \sqrt{\frac{\hat p (1 - \hat p ) }{ n} } } $$

OpenStudy (anonymous):

we're trying to find confidence mean?

OpenStudy (anonymous):

would we look for the confidence proportion since we don't have any knowledge as to what the population size is?

OpenStudy (perl):

it depends on the problem, what you are sampling. if you are sampling yes/no to a political survey , then you would want proportion. if you are sampling say mean height of freshman college students, you would use sample mean.

OpenStudy (anonymous):

the problem doesn't say :/

OpenStudy (perl):

i think this problem from the context you can tell they want a confidence interval for the sample mean

OpenStudy (perl):

since it says 'the mean is 20' , they mean the sample mean is 20

OpenStudy (anonymous):

okay, so how do we find that without population

OpenStudy (perl):

if they wanted proportion it would be different , it would say the sample proportion is say 60%

OpenStudy (anonymous):

how do we calculate this? I'm confused

OpenStudy (perl):

$$ \Large{ \text{confidence interval for sample mean is: } \\ \bar x \pm z^* \frac{\sigma }{\sqrt n} \\\therefore\\ 20 \pm 1.96 \cdot \frac{3 }{\sqrt {30}} } $$

OpenStudy (anonymous):

oh we would use 30 for the population??

OpenStudy (perl):

\(\large \sigma \) is the population standard deviation. n is the size of your sample

OpenStudy (anonymous):

20 +/- 1.073536

OpenStudy (perl):

\( \large \bar x \) is the sample mean (of the 30 data points)

OpenStudy (perl):

yes thats correct :)

OpenStudy (anonymous):

okay thank you so much!! i was so stuck

OpenStudy (anonymous):

you're the best

OpenStudy (anonymous):

wait sorry @perl is my margin of error of 5.88 or no?